OpenStudy (anonymous):

Find the slope of the line. (-2,3) and (5,3)

OpenStudy (ivettef365):

you need to use slope formula, do you know it ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

a.1/7 b.-1/7 c.-7 d.7

OpenStudy (ivettef365):

y2 - y1 ------ x2 - x1

OpenStudy (anonymous):

3-3=0 5 -( -2)=7

OpenStudy (anonymous):

so is D

OpenStudy (anonymous):

the answer

OpenStudy (ivettef365):

actually is not D, because you have 0/7 which = 0 slope = 0

OpenStudy (ivettef365):

are you sure you have the point written correctly

OpenStudy (anonymous):

yeah these what in the graph

OpenStudy (anonymous):

may they should be that way (5,3) and (-2,3)

OpenStudy (ivettef365):

the way does not alter the slope, still 0

OpenStudy (ivettef365):

you have 2 y's with the same number that is 0, when you have 2 x's with the same number then the slope is undefined

OpenStudy (ivettef365):

are you taking the points from a graph?

OpenStudy (anonymous):

yeap

OpenStudy (ivettef365):

does the graph have a straight line

OpenStudy (ivettef365):

or a tilted line?

OpenStudy (anonymous):

straight line

OpenStudy (ivettef365):

then those choices are incorrect

OpenStudy (anonymous):

oh ok thanks anyway